OpenStudy (anonymous):

What is the number of solutions and solution of 3x+2y=10 and 3x+2y=2 ??

OpenStudy (anonymous):

@Mertsj help please?

OpenStudy (anonymous):

they are identical on the left side, but different on the right what does that tell you?

OpenStudy (anonymous):

let us rewrite the equations in slope-intercept form. what do we get as slopes for the two lines?

OpenStudy (anonymous):

@electrokid the slope is 2

OpenStudy (anonymous):

\(y = mx+b\) re-arrange the equations like this.

OpenStudy (anonymous):

2y=3x+10 and 2y=3x+2

OpenStudy (anonymous):

yes. then divide by "2"

OpenStudy (anonymous):

because we want "y" on its own

OpenStudy (anonymous):

let me show the first one \[3x+2y=10\\ -3x\quad-3x\\ ------\\ 2y=-3x+10\\\text{divide by 2}\\ {2y\over 2}={-3\over 2}x+{10\over 2}\\ \boxed{y={-3\over2}x+5} \]similarly, for the second one

OpenStudy (anonymous):

i got y=3/2+5 and y=3/2+1

OpenStudy (anonymous):

you missed the "x" in them

OpenStudy (anonymous):

oh yeah my bad but i do know what you mean, and my next step would be what?

OpenStudy (anonymous):

what did you get for the second equation?

OpenStudy (anonymous):

and what do we get for the slopes of each line?

OpenStudy (anonymous):

well?

OpenStudy (anonymous):

i will butt in for a second, and then butt out how is it possible for \(3x+2y\) to equal two different numbers? it is not

OpenStudy (anonymous):

the slope is 3/2 for both

OpenStudy (anonymous):

-3/2 for both and lines with equal slopes are parallel to each other. since parallel lines never intersect, there are NO SOLUTIONS for the given equations

OpenStudy (anonymous):

kapeesh?

OpenStudy (anonymous):

kapeesh xD and my hw also asks for the solution though?

OpenStudy (anonymous):

ok i will but back in if \(3x+2y=10\) then it cannot also equal 2, and vice versa

OpenStudy (anonymous):

so, what will your answer be for that? @satellite73 that is logical. Teachers usually want the answer using the "textual" method :/

OpenStudy (anonymous):

idk i got 0

OpenStudy (anonymous):

@electrokid ^

OpenStudy (anonymous):

0 what? 0 number of solutions there are no solutions.. i.e., there is no such pair of (x,y) that will lie on both the lines since they are parallel 

OpenStudy (anonymous):

so then why would my teacher put solutiion in my hw if there isnt any?

OpenStudy (anonymous):

to trick ya

OpenStudy (anonymous):

so im just suppose to put 0,0 on there @electrokid

OpenStudy (anonymous):

(0,0) is a point. so you cannot put that. what is the number of solutions? __ what are the solutions? __ explain

OpenStudy (anonymous):

the solutions are 5 and 1

OpenStudy (anonymous):

@electrokid is it right?